Calculate the concentrations of all species in a 0.290 M Na2SO3 0.290 M Na2SO3 (sodium sulfite) solution. The ionization constants for sulfurous acid are

0 1
Answer #1

[Na+] = 2 × 0.290 M = 0.580M. (Answer)

[SO32-] = 0.290M

[HSO3-] = 2.146×10-4 M

[H2SO3]= 7.14×10-13M

[OH-]= 2.146×10-4

[H+] = 4.66×10-11M

6.3 158 [Wgse,าะ 7.lyX1o13m